# I Introduction

### v1

@@ Abstract
Optimizing the performance of large language models (LLMs) on large-scale AI training and inference systems requires a scalable and expressive mechanism to model distributed workload execution. Such modeling is essential for pre-deployment system-level optimizations (e.g., parallelization strategies) and hardware design-space explorations. While recent efforts have proposed collecting execution traces from real systems, access to large-scale infrastructure remains limited to major cloud providers. Moreover, traces capturing execution on a specific platform cannot be easily adapted to study alternate software and/or hardware configurations, especially at scale. We introduce STAGE ***Symbolic Tensor grAph GEnerator, a framework that synthesizes high-fidelity execution graphs to accurately model distributed AI workloads (including LLMs and MoEs). STAGE supports a comprehensive set of parallelization strategies, allowing users to systematically explore a wide spectrum of model architectures and system configurations. STAGE demonstrates its scalability by synthesizing high-fidelity LLM traces spanning over 32K GPUs, while preserving tensor-level accuracy in compute, memory, and communication. STAGE will be publicly available to facilitate further research in distributed machine learning systems.
11footnotetext: Equal contribution.
@@ I Introduction
  • Validation with Real-World Traces: STAGE generates execution graphs that model computation, communication, and memory behavior, and we validate their fidelity using real traces collected from a single GPU to a production-scale 128-GPU H100/H200 HGX clusters executing large-scale LLM training workloadsUpon acceptance of this paper, we will release these traces as publicly available complementary resources.
@@ V-B Memory Footprint Validation
For memory‑footprint validation, we fed STAGE ‑synthesized graphs to ASTRA‑Sim [astraSim], extending its native Chakra‑format workload feederASTRA‑Sim natively consumes execution traces in the Chakra format. to track memory usage over the simulation lifetime. Our modifications enable ASTRA‑Sim to utilize tensor metadata (e.g., name, size) from STAGE graphs when generating tensor read/write events. These events are then post‑processed to determine each tensor’s lifetime, from creation to last use, assuming garbage collection immediately thereafter.
@@ VI-B Scalability Studies with STAGE
Target System Setup: The simulated system is constructed using NVIDIA DGX systems, where each box contains 8 H100 GPUs interconnected via NVLink. Sixteen boxes form a pod interconnected through a local ring, and multiple pods are connected via a ring topology. Our experiments span system sizes from 512 GPUs up to 16K GPUs§§§Because we are targeting the large-scale systems, the native ASTRA-sim simulation triggers an Out of Memory failure. We enhanced the workload feeder for offloaded trace processing on hard drive with caching.

### v2

@@ Abstract
Optimizing the performance of large language models (LLMs) on large-scale AI training and inference systems requires a scalable and expressive mechanism to model distributed workload execution. Such modeling is essential for pre-deployment system-level optimizations (e.g., parallelization strategies) and hardware design-space explorations. While recent efforts have proposed collecting execution traces from real systems, access to large-scale infrastructure remains limited to major cloud providers. Moreover, traces capturing execution on a specific platform cannot be easily adapted to study alternate software and/or hardware configurations, especially at scale. We introduce STAGE  [^1] Symbolic Tensor grAph GEnerator, a framework that synthesizes high-fidelity execution graphs to accurately model distributed AI workloads (including LLMs and MoEs). STAGE supports a comprehensive set of parallelization strategies, allowing users to systematically explore a wide spectrum of model architectures and system configurations. STAGE demonstrates its scalability by synthesizing high-fidelity LLM traces spanning over 32K GPUs, while preserving tensor-level accuracy in compute, memory, and communication. STAGE is publicly available to facilitate further research in distributed machine learning systems: https://github.com/astra-sim/symbolic_tensor_graph
@@ I Introduction
  • Validation with Real-World Traces: STAGE generates execution graphs that model computation, communication, and memory behavior, and we validate their fidelity using real traces collected from a single GPU to a production-scale 128-GPU H100/H200 HGX clusters executing large-scale LLM training workloads
@@ V-B Memory Footprint Validation
For memory‑footprint validation, we fed STAGE ‑synthesized graphs to ASTRA‑Sim [astraSim], extending its native Chakra‑format workload feeder [^2] ASTRA‑Sim natively consumes execution traces in the Chakra format. to track memory usage over the simulation lifetime. Our modifications enable ASTRA‑Sim to utilize tensor metadata (e.g., name, size) from STAGE graphs when generating tensor read/write events. These events are then post‑processed to determine each tensor’s lifetime, from creation to last use, assuming garbage collection immediately thereafter.
@@ VI-B Scalability Studies with STAGE
Target System Setup: The simulated system is constructed using NVIDIA DGX systems, where each box contains 8 H100 GPUs interconnected via NVLink. Sixteen boxes form a pod interconnected through a local ring, and multiple pods are connected via a ring topology. Our experiments span system sizes from 512 GPUs up to 16K GPUs [^3] Because we are targeting the large-scale systems, the native ASTRA-sim simulation triggers an Out of Memory failure. We enhanced the workload feeder for offloaded trace processing on hard drive with caching.
@@ VIII Conclusion
Acknowledgments
This work was supported, in part, by ACE, one of the seven centers in JUMP2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. We thank Michael Weiner and Eric Coulter from the PACE team at Georgia Tech, and Matthieu Bloch, for provisioning access to the Georgia Tech AI Makerspace for Chakra trace collection. We also appreciate early feedback for this work from the Chakra Working Group at MLCommons.